Write the equation of the line, in a slope-intercept form, that passes through the points (5, 3) and (4,10).
Answer by subudear(62)

Line equation is given by y = mx +c
m = (y2 - y1)/(x2 - x1)
lets say (x1, y1) = (5,3) and (x2,y2) = (4,10)
then m = (10 - 3)/(4 - 5) = -7
equation of line is
y - 3 = (-7)(x -5)
y - 3 = -7x +35
y = -7x + 38